the perimeter of a rectangle is 64 feet..the width is 1 foot shorter than 1/2 the length..
i thought the equation was : l + 1/2-1 =64...not coming up with right answers.

Answer by rothauserc(4718)

let x be the length, then width = (x/2) - 1
Perimeter(P) = 2w + 2l
P = 2((x/2) -1) + 2x
P = 2x/2 -2 + 2x
P = x - 2 +2x = 3x - 2
P is 64
64 = 3x -2
3x = 66
x = 22
therefore
length is 22 feet and
width = (22/2) - 1
width = 10
length is 22 feet and width is 10 feet